### 28 Nov 2023 v1

# Photo-SLAM: Real-time Simultaneous Localization and Photorealistic Mapping for Monocular, Stereo, and RGB-D Cameras

Abstract
The integration of neural rendering and the SLAM system recently showed promising results in joint localization and photorealistic view reconstruction. However, existing methods, fully relying on implicit representations, are so resource-hungry that they cannot run on portable devices, which deviates from the original intention of SLAM. In this paper, we present Photo-SLAM, a novel SLAM framework with a hyper primitives map. Specifically, we simultaneously exploit explicit geometric features for localization and learn implicit photometric features to represent the texture information of the observed environment. In addition to actively densifying hyper primitives based on geometric features, we further introduce a Gaussian-Pyramid-based training method to progressively learn multi-level features, enhancing photorealistic mapping performance. The extensive experiments with monocular, stereo, and RGB-D datasets prove that our proposed system Photo-SLAM significantly outperforms current state-of-the-art SLAM systems for online photorealistic mapping, e.g., PSNR is 30% higher and rendering speed is hundreds of times faster in the Replica dataset. Moreover, the Photo-SLAM can run at real-time speed using an embedded platform such as Jetson AGX Orin, showing the potential of robotics applications.
1 Introduction
Figure 1: Photo-SLAM is a novel framework for simultaneous localization and photorealistic mapping in real-time supporting monocular, stereo, and RGB-D cameras. It can reconstruct high-fidelity views of scenes while render speed is up to 1000 FPS.
Simultaneous Localization and Mapping (SLAM) using cameras is a fundamental problem in both computer vision and robotics, seeking to enable autonomous systems to navigate and comprehend their surroundings. Traditional SLAM systems [24, 9, 7, 8] primarily focus on geometric mapping, providing accurate but visually simplistic representations of the environment. However, recent developments in neural rendering [34, 39] have demonstrated the potential of integrating photorealistic view reconstruction into the SLAM pipeline, enhancing the perception capabilities of robotic systems.
Despite the promising results achieved through the integration of neural rendering and SLAM, existing methods simply and heavily rely on implicit representations, making them computationally intensive and unsuitable for deployment on resource-constrained devices. For example, Nice-SLAM [45] leverages a hierarchical grid [41] to store learnable features representing the environment while ESLAM [16] utilizes multi-scale compact tensor components [3]. They then jointly estimate the camera poses and optimize features by minimizing the reconstruction loss of a batch of ray sampling [21]. Such an optimization process is time-consuming. Consequently, it is indispensable for them to incorporate corresponding depth information obtained from various sources such as RGB-D cameras, dense optical flow estimators [32], or monocular depth estimators [12] to ensure efficient convergence. Additionally, since the implicit features are decoded by the multi-layer perceptrons (MLP), it is typically necessary to carefully define a bounding area to normalize ray sampling for optimal performance [14]. It essentially limits the scalability of the system. These limitations imply that they cannot provide real-time exploration and mapping capabilities in the unknown environment using portable platforms, which is one of the main objectives of SLAM.
In this paper, we propose Photo-SLAM, an innovative framework that addresses the scalability and computational resource constraints of existing methods, while achieving precise localization and online photorealistic mapping. We maintain a hyper primitives map which is composed of point clouds storing ORB features [26], rotation, scaling, density, and spherical harmonic (SH) coefficients [37, 10]. The hyper primitives map allows the system to efficiently optimize tracking using a factor graph solver and learn the corresponding mapping by backpropagating the loss between the original images and rendering images. The images are rendered by 3D Gaussian splatting [18] rather than ray sampling. Although the introduction of a 3D Gaussian splatting renderer can reduce view reconstruction costs, it does not enable the generation of high-fidelity rendering for online incremental mapping, in particular in monocular scenarios. To achieve high-quality mapping without reliance on dense depth information, we further propose a geometry-based densification strategy and a Gaussian-Pyramid-based (GP) learning method. Importantly, GP learning facilitates the progressive acquisition of multi-level features which effectively enhances the mapping performance of our system.
To evaluate the efficacy of our proposed approach, we conduct extensive experiments employing diverse datasets captured by monocular, stereo, and RGB-D cameras. These experiment results unequivocally demonstrate that Photo-SLAM attains state-of-the-art performance in terms of localization efficiency, photorealistic mapping quality, and rendering speed. Furthermore, the real-time execution of the Photo-SLAM system on the embedded devices showcases its potential for practical robotics applications. The schematic overview of Photo-SLAM is demonstrated in Fig. 1 and Fig. 1(b).
In summary, the main contributions of this work include:
  • We developed the first simultaneous localization and photorealistic mapping system based on hyper primitives map. The novel framework supports monocular, stereo, and RGB-D cameras, in indoor and outdoor environments.
  • We proposed Gaussian-Pyramid-based learning allowing the model to efficiently and effectively learn multi-level features realizing high-fidelity mapping.
  • The system, fully implemented in C++ and CUDA, achieves start-of-the-art performance and can run at real-time speed even on embedded platforms. The code will be publicly available.
2 Related Work
Figure 2: The Photo-SLAM contains four main components, including localization, explicit geometry mapping, implicit photorealistic mapping, and loop closure components, while maintaining a map with hyper primitives.
Visual localization and mapping is a problem that aims to build a proper representation of an unknown environment via cameras while estimating their poses within that environment. In contrast to SfM techniques, visual SLAM techniques typically pursue a better trade-off between accuracy and real-time performance. In this section, we focus on visual SLAM and conduct a brief review.
Graph Solver vs Neural Solver. Classical SLAM methods widely adopt factor graphs to model complex optimization problems between variables (i.e., poses and landmarks) and measurements (i.e., observations and constraints). To achieve real-time performance, SLAM methods incrementally propagate their pose estimations while avoiding expensive operations. For example, ORB-SLAM series methods  [24, 23, 2] rely on extracting and tracking lightweight geometric features across consecutive frames, which perform bundle adjustment locally instead of globally. Moreover, direct SLAMs like LSD-SLAM [7] and DSO [8] operate on raw image intensities, without the cost of geometric feature extractions. They maintain a sparse or semi-dense map represented by point clouds online, even on the resource-constraint system. Benefiting from the success of deep-learning models, learnable parameters and models are introduced into SLAM making the pipeline differentiable. Some methods such as DeepTAM [44] predict camera poses by the neural network [17] end-to-end, while the accuracy is limited. To enhance performance, some methods, e.g., D3VO [40] and Droid-SLAM  [33], introduce monocular depth estimation [12] or dense optical flow estimation [32] models into the SLAM pipeline as supervision signals. Therefore, they can generate depth maps that explicitly represent the scene geometry. With the large-scale synthetic SLAM dataset, TartanAir [36], available for training, Droid-SLAM building upon RAFT [32] achieves state-of-the-art performance. However, the pure neural-based solver is computationally expensive and their performance would significantly degrade on the unseen scenes.
Explicit Representation vs Implicit Representation. In order to obtain dense reconstruction, some methods including KinectFusion [15], BundleFusion [6], and InfiniTAM [25] utilize the implicit representation, Truncated Signed Distance Function (TSDF) [5], to integrate the incoming RGB-D images and reconstruct a continuous surface, which can run in real time on GPU. Although they can obtain dense reconstruction, view rendering quality is limited. Recently, neural rendering techniques represented by neural radiance field (NeRF) [21] have achieved breathtaking novel view synthesis. Given camera poses, NeRF implicitly models the scene geometry and color by multi-layer perceptrons (MLP). The MLP is optimized by minimizing the loss of rendering images and training views. iMAP [29] then adapts NeRF for incremental mapping, optimizing not only MLP but also camera poses. The following work Nice-SLAM [45] introduces multi-resolution grids [41] to store features reducing the cost of deep MLP query. Co-SLAM [35] and ESLAM [16] explore Instant-NGP [22] and TensoRF [3] respectively to further accelerate the mapping speed. However, implicitly joint optimization of camera poses and geometry representation is still ill-conditioned. Inevitably, they rely on explicit depth information from RGB-D cameras or additional model predictions for fast convergence of the radiance field.
Our proposed Photo-SLAM seeks to recover a concise representation of the observed environment for immersive exploration rather than reconstructing a dense mesh. It maintains a map with hyper primitives online which capitalizes on explicit geometric feature points for accurate and efficient localization while leveraging implicit representations to capture and model the texture information. Please refer to Fig. 1(a) for the taxonomy of existing systems. Since Photo-SLAM achieves high-quality mapping without reliance on dense depth information, it can support RGB-D cameras as well as monocular and stereo cameras.
3 Photo-SLAM
Photo-SLAM contains four main components, including localization, geometry mapping, photorealistic mapping, and loop closure, shown in Fig. 1(b). Each component runs in a parallel thread and jointly maintains a hyper primitives map.
Figure 3: Comparison of different progressive training methods. The encoder $\mathcal{E}_{n}$ here represents a structure to regress features $\mathcal{F}_{n}$ which can be an MLP, voxel grid, hash table, positional encoding, etc. The decoder $\mathcal{D}_{n}$ here represents a structure converting $\mathcal{F}_{n}$ into density, color, or other information. We proposed a new method based on the Gaussian pyramid to efficiently learn multi-level features.
3.1 Hyper Primitives Map
In our system, hyper primitives are defined as a set of point clouds $\mathbf{P}\in\mathbb{R}^{3}$ associated with ORB features [26] $\mathbf{O}\in\mathbb{R}^{256}$, rotation $\mathbf{r}\in SO(3)$, scaling $\mathbf{s}\in\mathbb{R}^{3}$, density $\sigma\in\mathbb{R}^{1}$, and spherical harmonic coefficients $\mathbf{SH}\in\mathbb{R}^{16}$. ORB features extracted from image frames take responsibility for establishing 2D-to-2D and 2D-to-3D correspondences. Once the system successfully estimates the transformation matrix based on sufficient 2D-to-2D correspondences between adjacent frames, the hyper primitives map is initialized via triangulation, and pose tracking gets started. During tracking, the localization component processes the incoming images and makes use of 2D-to-3D correspondence to calculate current camera poses. In addition, the geometry mapping component will incrementally create and initialize sparse hyper primitives. Finally, the photorealistic component progressively optimizes and densifies hyper primitives.
3.2 Localization and Geometry Mapping
The localization and geometry mapping components provide not only efficient 6-DoF camera pose estimations of the input images, but also sparse 3D points. The optimization problem is formulated as a factor graph solved by the Levenberg–Marquardt (LM) algorithm.
In the localization thread, we use a motion-only bundle adjustment to optimize the camera orientation $\mathbf{R}\in SO(3)$ and position $\mathbf{t}\in\mathbb{R}^{3}$ in order to minimize the reprojection error between matched 2D geometric keypoint $\mathbf{p}_{i}$ of the frame and 3D point $\mathbf{P}_{i}$. Let $i\in\mathcal{X}$ be the index of set of matches $\mathcal{X}$, what we are trying to optimize with LM is
$\{\mathbf{R},\mathbf{t}\}=\mathop{\arg\min}\limits_{\mathbf{R},\mathbf{t}}\sum_{i\in\mathcal{X}}\rho\left(\lVert\mathbf{p}_{i}-\pi(\mathbf{RP}_{i}+\mathbf{t})\rVert^{2}_{\Sigma_{g}}\right),$ (1)
where $\Sigma_{g}$ is the scale-associated covariance matrix of the keypoint, $\pi(\cdot)$ is the 3D-to-2D projection function, and $\rho$ denotes the robust Huber cost function.
In the geometry mapping thread, we perform a local bundle adjustment on a set of covisible points $\mathcal{P}_{L}$ and keyframes $\mathcal{K}_{L}$. The keyframes are selected frames from the input camera sequence and provide good visual information. We construct a factor graph where each keyframe is a node, and the edges represent constraints between keyframes and matched 3D points. We iteratively minimize the reprojection residual by refining the keyframe poses and 3D points using the first-order derivatives of the error function. We fix the poses of keyframes $\mathcal{K}_{F}$ which are also observing $\mathcal{P}_{L}$ but not in $\mathcal{K}_{L}$. Let $\mathcal{K}=\mathcal{K}_{L}\cup\mathcal{K}_{F}$, and $\mathcal{X}_{k}$ be the set of matches between 2D keypoints in a keyframe $k$ and 3D points in $\mathcal{P}_{L}$. The optimization process aims to reduce the geometric inconsistency between $\mathcal{K}$ and $\mathcal{P}_{L}$, and is defined as
$\{\mathbf{P}_{i},\mathbf{R}_{l},\mathbf{t}_{l}|i\in\mathcal{P}_{L},l\in\mathcal{K}_{L}\}=\mathop{\arg\min}\limits_{\mathbf{P}_{i},\mathbf{R}_{l},\mathbf{t}_{l}}\sum_{k\in\mathcal{K}}\sum_{j\in\mathcal{X}_{k}}\rho(E(k,j)),$ (2)
with reprojection residual
$E(k,j)=\lVert\mathbf{p}_{j}-\pi(\mathbf{R}_{k}\mathbf{P}_{j}+\mathbf{t}_{k})\rVert^{2}_{\Sigma_{g}}.$
3.3 Photorealisitc Mapping
The photorealistic mapping thread is responsible for optimizing hyper primitives that are incrementally created by the geometry mapping thread. The hyper primitives can be rasterized by a tile-based renderer to synthesize corresponding images with keyframe poses. The rendering process is formulated as
$C(\mathbf{R,t})=\sum_{i\in N}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}),$ (3)
where $N$ is the number of hyper primitives, $\mathbf{c}_{i}$ denotes the color converted from $\mathbf{SH}\in\mathbb{R}^{16}$, and $\alpha_{i}$ is equal to $\sigma_{i}\cdot\mathcal{G}(\mathbf{R,t},\mathbf{P}_{i},\mathbf{r}_{i},\mathbf{s}_{i})$, $\mathcal{G}$ denotes 3D Gaussian splatting algorithm [18]. The optimization in terms of position $\mathbf{P}$, rotation $\mathbf{r}$, scaling $\mathbf{s}$, density $\sigma$, and spherical harmonic coefficients $\mathbf{SH}$ is performed by minimizing the photometric loss $\mathcal{L}$ between rendering image ${I_{\text{r}}}$ and ground truth image $I_{\text{gt}}$, denoted as
$\mathcal{L}=(1-\lambda)\left|{I_{\text{r}}-I_{\text{gt}}}\right|_{1}+\lambda\frac{1-\text{SSIM}(I_{\text{r}},I_{\text{gt}})}{2},$ (4)
where $\text{SSIM}(I_{\text{r}},I_{\text{gt}})$ denotes structural similarity between two images and $\lambda$ is a weight factor for balance.
3.3.1 Geometry-based Densification
If we consider photorealistic mapping as a regression model of the scene, denser hyper primitives, i.e., more parameters, generally can better model the complexity of the scene for higher rendering quality. To meet the demand for real-time mapping, the geometry mapping component only establishes sparse hyper primitives. Therefore, the coarse hyper primitives created by the geometry mapping need to be densified during the optimization of photorealistic mapping. Apart from splitting or cloning hyper primitives with large loss gradients similar to [18], we introduce an additional geometry-based densification strategy.
Experimentally, less than $30\%$ of 2D geometric feature points of frames are active and have corresponding 3D points, especially for non-RGB-D scenarios, as shown in Fig. 4. We argue that 2D geometric feature points spatially distributed in the frames essentially represent the region with a complex texture that requires more hyper primitives. Therefore, we actively create additional temporary hyper primitives based on the inactive 2D feature points once the keyframe is created for photorealistic mapping. When we use RGB-D cameras, we can directly project the inactive 2D feature points with depth to create temporary hyper primitives. As for monocular scenarios, we estimate the depth of inactive 2D feature points by interpreting the depth of their nearest neighborhood’s active 2D feature points. In stereo scenarios, we rely on a stereo-matching algorithm to estimate the depth of inactive 2D feature points.
Figure 4: We make use of initial geometric information to densify hyper primitives.
3.4 Gaussian-Pyramid-Based Learning
Progressive training is a widely used technology in neural rendering to accelerate the optimization process. Some methods have been proposed to reduce training time while achieving better rendering quality. A basic method is to progressively increase the structure resolution and the number of model parameters. For example, NSVF [20] and DVGO [30] progressively increase the feature grid resolution during training which significantly improves training efficiency compared to previous work. The lower-resolution model is used to initialize the higher-resolution model but is not retained for final inference, as shown in Fig. 2(a). To enhance performance with multi-resolution features, NGLoD [31] progressively trains multiple MLPs as encoders and decoders, while only retaining the final decoder to decode integrated multi-resolution features, as shown in Fig. 2(b). Furthermore, Neuralangelo [19] only maintains a single MLP during training, as shown in Fig. 2(c). It progressively activates different levels of hash tables [22] achieving better performance in large-scale scene reconstruction. Similarly, 3D Gaussian Splatting [18] progressively densifies 3D Gaussian achieving top performance on radiance field rendering. Training different level models in these methods is supervised by the same training images. Conversely, the fourth method (Fig. 2(d)) used in BungeeNeRF [38] is to apply different models to tackle different-resolution images. BungeeNeRF demonstrates the efficiency of explicitly grouping multi-resolution training images for models to learn multi-level features. However, such a method is not universal since multi-resolution images are not available for most scenarios.
To make full use of various merits, we propose Gaussian-Pyramid-based (GP) learning (Fig. 2(e)), a new progressive training method. As illustrated in Fig. 5, a Gaussian pyramid is a multi-scale representation of an image containing different levels of detail. It is constructed by repeatedly applying Gaussian smoothing and downsampling operations to the original image. At the beginning training step, the hyper primitives are supervised by the highest level of the pyramid, i.e. level $n$. As training iteration increases, we not only densify hyper primitives as described in Sec. 3.3.1 but also reduce the pyramid level and obtain a new ground truth until reaching the bottom of the Gaussian pyramid. The optimization process using a Gaussian pyramid with n+1 levels can be denoted as
$\begin{split}t_{0}&:\mathop{\arg\min}\mathcal{L}\left(I^{n}_{\text{r}},\text{GP}^{n}(I_{\text{gt}})\right),\\ t_{1}&:\mathop{\arg\min}\mathcal{L}\left(I^{n-1}_{\text{r}},\text{GP}^{n-1}(I_{\text{gt}})\right),\\ &\dots\\ t_{n}&:\mathop{\arg\min}\mathcal{L}\left(I^{0}_{\text{r}},\text{GP}^{0}(I_{\text{gt}})\right),\end{split}$ (5)
where $\mathcal{L}(I_{\text{r}},\text{GP}(I_{\text{gt}}))$ is Eq. 4, while $\text{GP}^{n}(I_{\text{gt}})$ denotes the ground image in the level n of the Gaussian pyramid. In the experiment, we prove that GP learning significantly improves the performance of photorealistic mapping particularly for monocular cameras.
Figure 5: Training process based on the Gaussian pyramid.
3.5 Loop Closure
Loop Closure [11] is crucial in SLAM because it helps address the problem of accumulated errors and drift that can occur during the localization and geometry mapping process. After detecting a closing loop, we can correct local keyframes and hyper primitives by similarity transformation. With corrected camera poses, the photorealistic mapping component can further get rid of the ghosting caused by odometry drifts and improve the mapping quality.
4 Experiment
In this section, we compare Photo-SLAM to other state-of-the-art (SOTA) SLAM and real-time 3D reconstruction systems in various scenarios encapsulating monocular, stereo, RGB-D cameras, and indoor and outdoor environments. In addition, we evaluate Photo-SLAM performance on various hardware configurations to demonstrate its efficiency. Finally, we conduct an ablation study to verify the effectiveness of the proposed algorithms.
On Replica Dataset Localization (cm) Mapping Resources
Cam Method RMSE $\downarrow$ STD $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Tracking FPS $\uparrow$ Rendering FPS $\uparrow$ GPU Memory Usage $\downarrow$
Mono ORB-SLAM3 [2] 3.942 3.115 - - - 58.749 - 0
DROID-SLAM [33] 0.725 0.308 - - - 35.473 - 11 GB
Nice-SLAM* [45] 99.9415 35.336 16.311 0.720 0.439 2.384 0.944 12 GB
Orbeez-SLAM [4] - - 23.246 0.790 0.336 49.200 1.030 6 GB
Go-SLAM [43] 71.054 24.593 21.172 0.703 0.421 25.366 0.821 22 GB
Ours (Jetson) 1.235 0.756 29.284 0.883 0.139 18.315 95.057 4 GB
Ours (Laptop) 0.713 0.524 33.049 0.926 0.086 19.974 353.504 4 GB
Ours 1.091 0.892 33.302 0.926 0.078 41.646 911.262 6 GB
RGB-D ORB-SLAM3 [2] 1.833 1.478 - - - 52.209 - 0
DROID-SLAM [33] 0.634 0.248 - - - 36.452 - 11 GB
BundleFusion [6] 1.606 0.969 23.839 0.822 0.197 8.630 - 5 GB
Nice-SLAM [45] 2.350 1.590 26.158 0.832 0.232 2.331 0.611 12 GB
Orbeez-SLAM [4] 0.888 0.562 32.516 0.916 0.112 41.333 1.401 6 GB
ESLAM [16] 0.568 0.274 30.594 0.866 0.162 6.687 2.626 21 GB
Co-SLAM [35] 1.158 0.602 30.246 0.864 0.175 14.575 3.745 4 GB
Go-SLAM [43] 0.571 0.218 24.158 0.766 0.352 19.437 0.444 24 GB
Ours (Jetson) 0.581 0.289 31.978 0.916 0.101 17.926 116.395 4 GB
Ours (Laptop) 0.590 0.289 34.853 0.944 0.062 20.597 396.082 4 GB
Ours 0.604 0.298 34.958 0.942 0.059 42.485 1084.017 5 GB
Table 1: Quantitative results on the Replica dataset. We mark the best two results with first and second. Nice-SLAM* means the depth supervision is disabled. “-” denotes the system does not support view rendering or fails to track camera poses. The results of Photo-SLAM running on the laptop and Jetson platform are denoted as “Ours (Laptop)” and “Ours (Jetson)” respectively.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 6: Qualitative comparison of diverse systems using RGB-D images from dataset Replica. Photo-SLAM can reconstruct high-fidelity scenes while others are over-smooth and have obvious artifacts. Zoom in for better views.
Figure 7: Mapping comparison with other monocular camera systems.
On TUM Dataset fr1-desk fr2-xyz fr3-office
Cam Method RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Mono ORB-SLAM3 [2] 1.534 - - - 0.720 - - - 1.400 - - -
DROID-SLAM [33] 78.245 - - - 36.050 - - - 154.383 - - -
Go-SLAM [43] 33.122 11.705 0.406 0.614 28.584 14.807 0.443 0.572 105.755 13.572 0.480 0.643
Ours (Jetson) 1.757 18.811 0.681 0.329 0.558 21.347 0.727 0.187 1.687 18.884 0.672 0.289
Ours (Laptop) 1.549 20.515 0.733 0.241 0.852 21.575 0.739 0.157 1.542 19.138 0.680 0.259
Ours 1.539 20.972 0.743 0.228 0.984 21.072 0.726 0.166 1.257 19.591 0.692 0.239
RGB-D ORB-SLAM3 [2] 1.724 - - - 0.385 - - - 1.698 - - -
DROID-SLAM [33] 91.985 - - - 41.833 - - - 160.141 - - -
Nice-SLAM [45] 19.317 12.003 0.417 0.510 36.103 18.200 0.603 0.313 25.309 16.341 0.548 0.386
ESLAM [16] 3.359 17.497 0.561 0.484 31.448 22.225 0.727 0.233 25.808 19.113 0.616 0.359
Co-SLAM [35] 3.094 16.419 0.482 0.591 31.347 19.176 0.595 0.374 25.374 17.863 0.547 0.452
Go-SLAM [43] 2.119 15.794 0.531 0.538 31.788 16.118 0.534 0.419 26.802 16.499 0.566 0.569
Ours (Jetson) 4.571 18.273 0.663 0.338 0.360 23.127 0.780 0.149 1.874 19.781 0.701 0.235
Ours (Laptop) 1.891 20.403 0.728 0.251 0.361 22.570 0.777 0.158 1.315 21.569 0.749 0.184
Ours 2.603 20.870 0.743 0.239 0.346 22.094 0.765 0.169 1.001 22.744 0.780 0.154
Table 2: Quantitative results on the TUM RGB-D dataset. We mark the best two results with first and second.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 8: Qualitative results of Photo-SLAM on dataset TUM.
4.1 Implementation and Experiment Setup
We implemented Photo-SLAM fully in C++ and CUDA, making use of ORB-SLAM3 [2], 3D Gaussian splatting [18], and the LibTorch framework. The optimization of photorealistic mapping is performed with the Stochastic Gradient Descent algorithm while we use a fixed learning rate and $\lambda=0.2$. Considering image resolution of the testing dataset, the level of the Gaussian pyramid is set to three, i.e. $n=2$ by default. The compared baseline includes a SOTA classical SLAM system ORB-SLAM3 [2], a real-time RGB-D dense reconstruction system BundleFusion [6], a deep-learning based system DROID-SLAM [33], and recent SLAM systems supporting view synthesis, i.e. Nice-SLAM [45], Orbeez-SLAM [4], ESLAM [16], Co-SLAM [35], and Go-SLAM [43].
Hardware. We ran Photo-SLAM and all compared methods using their official code in a desktop with an NVIDIA RTX 4090 24 GB GPU, an Intel Core i9-13900K CPU, and 64 GB RAM. We further tested Photo-SLAM on a laptop and a Jetson AGX Orin Developer Kit. The laptop is equipped with an NVIDIA RTX 3080ti 16 GB Laptop GPU, an Intel Core i9-12900HX, and 32 GB RAM.
Datasets and Metrics. We performed tests for monocular and RGB-D sensor types on the well-known RGB-D datasets: the Replica dataset [27, 29] and the TUM RGB-D dataset [28]. As for stereo tests, we used the EuRoC MAV dataset [1]. Besides indoor scenes, we utilize a ZED 2 stereo camera to collect outdoor scenes for extra evaluation.
Following the convention, we used the Absolute Trajectory Error (ATE) metric [13] to estimate the accuracy of localization, while the RMSE and STD of ATE are reported. Quantitative measurements in terms of PSNR, SSIM, and LPIPS [42] are adopted to analyze the performance of photorealistic mapping. We also report the requirement of computing resources by showing the tracking FPS, rendering FPS, and GPU memory usage. The evaluation regarding mesh reconstruction is out of the range of this work. Moreover, to lower the effect of the nondeterministic nature of multi-threading and machine-learning systems, we ran each sequence five times and reported the average results for each metric. Please refer to the supplementary for details.
On Euroc Stereo ORB- SLAM3 DROID- SLAM Ours (Jetson) Ours (Laptop) Ours
MH-01 RMSE (cm) $\downarrow$ 4.379 39.514 4.207 4.049 4.109
PSNR $\uparrow$ - - 13.979 13.962 13.952
SSIM $\uparrow$ - - 0.426 0.421 0.420
LPIPS $\downarrow$ - - 0.428 0.378 0.366
MH-02 RMSE (cm) $\downarrow$ 4.525 39.265 4.193 4.731 4.441
PSNR $\uparrow$ - - 14.210 14.254 14.201
SSIM $\uparrow$ - - 0.436 0.436 0.430
LPIPS $\downarrow$ - - 0.447 0.373 0.356
V1-01 RMSE (cm) $\downarrow$ 8.940 21.646 8.830 8.836 8.821
PSNR $\uparrow$ - - 16.933 17.025 17.069
SSIM $\uparrow$ - - 0.626 0.622 0.618
LPIPS $\downarrow$ - - 0.321 0.284 0.266
V2-01 RMSE (cm) $\downarrow$ 26.904 15.344 26.643 26.736 26.609
PSNR $\uparrow$ - - 16.038 16.052 15.677
SSIM $\uparrow$ - - 0.643 0.635 0.622
LPIPS $\downarrow$ - - 0.347 0.314 0.323
Table 3: Quantitative results on the EuRoC MAV dataset, using stereo inputs. Our Photo-SLAM is the first system to support online photorealistic mapping with stereo cameras.
Figure 9: Mapping results of Photo-SLAM using a hand-held stereo camera in an outdoor unbounding scene.
4.2 Results and Evaluation
On Replica. As quantitative comparison demonstrated in Table 1, Photo-SLAM achieves top performance in terms of mapping quality. With competitive localization accuracy, Photo-SLAM can track the camera poses in real time. Moreover, Photo-SLAM renders hundreds of photorealistic views in a resolution of 1200$\times$680 per second with less GPU memory usage. Even on the embedded platform, the rendering speed of Photo-SLAM is about 100 FPS.
In monocular scenarios, Photo-SLAM significantly suppresses other methods. When we disabled the depth supervision of Nice-SLAM [45], its accuracy of localization dramatically decreased while the mapping was of 16.311 PSNR. We conduct a qualitative comparison in Fig. 7. The mapping results of Photo-SLAM are photorealistic.
In RGB-D scenarios, we ran BundleFusion [6] with RGB-D sequences and then extracted textured mesh. And then we used a mesh render to render corresponding images for comparison. As shown in Fig. 6, the mesh reconstructed by the classical method is likely to be aliasing and hollow. ESLAM [16] and Go-SLAM [43] have the best localization accuracy, but the mapping lacks high-frequency details. By contrast, Photo-SLAM can render high-fidelity images and the rendering speed is about three hundred times faster.
On TUM. We provide quantitative analyses on the three sequences of the TUM dataset in Table 2. Compared to learning-based methods, e.g., DROID-SLAM [33] and Go-SLAM [43], ORB-SLAM3 runs faster without the requirement of GPU and has higher accuracy regarding localization. It is shown that the classical method still has advantages in terms of robustness and generalization. Fig. 8 is a gallery of Photo-SLAM mapping.
On Stereo. Stereo cameras can provide more robust tracking but have hardly been supported by former real-time dense SLAM systems. However, Photo-SLAM has been designed to be compatible with stereo cameras. We provide quantitative results on the EuRoC dataset in Table 3 and qualitative results in supplementary. The results show that our system could still perform decently in stereo scenes. Further, we used a hand-held stereo camera to collect some outdoor scenes, and the mapping results of Photo-SLAM are illustrated in Fig. 9.
On Replica Mono RGB-D
# Geo GP PSNR $\uparrow$ FPS $\uparrow$ MB PSNR $\uparrow$ FPS $\uparrow$ MB
(1) w/o n = 2 31.274 994.2 10.742 33.296 923.0 18.199
(2) w/ w/o 20.002 353.2 44.100 33.696 860.0 31.856
(3) w/o w/o 22.913 645.0 5.782 32.551 1010.8 13.901
(4) w/ n = 1 30.903 803.8 21.819 34.634 953.7 31.552
(5) w/ n = 3 31.563 877.6 22.510 33.305 946.2 31.039
default w/ n = 2 33.302 911.3 31.419 34.958 1084.0 35.211
Table 4: Ablation study on the effect of geometry-based densification (Geo) and Gaussian-Pyramid-based (GP) learning.
Figure 10: Mapping results of different ablated systems on the monocular Replica scene.
4.3 Ablation Study
We proposed geometry-based densification (Geo) and Gaussian-Pyramid-based (GP) learning to boost the system performance of real-time photorealistic mapping. In this section, we constructed an ablation study to measure the efficacy of each algorithm, which can be quantified by PSNR, rendering speed (FPS), and final model size in megabytes (MB). The quantitative results are demonstrated in Table 4.
The relationship is tangled between rendering quality or speed, and the number of hyper primitives represented by the model size. The models are typically small without actively densifying hyper primitives based on the inactive 2D geometric features (Geo). However, without Geo, PSNR suffers degradation by 2.028 and 1.662 on monocular and RGB-D scenarios respectively, as indicated in Table 4(1). Compared to Fig. 6(c), the rendering image without Geo (Fig. 9(a)) exhibits artifacts, such as on the ceiling. In RGB-D scenarios, more hyper primitives can get higher PSNR. However, without Gaussian-Pyramid-based learning, more hyper primitives are densified by Geo, and thus lead to a decrease in mapping quality and rendering speed especially in monocular scenarios, as visualized in Fig. 9(b) and reported in Table 4(2) and (3). This is because densified hyper primitives without precise depth information have inaccurate positions. Without thorough optimization, inaccurate hyper primitives become encumbrances. It is noticeable that the systems with GP learning can generally perform better, highlighting the effectiveness of GP learning. Additionally, increasing the Gaussian pyramid levels improves mapping quality, and our Photo-SLAM with default 3-level GP learning achieves the best results. However, we found that the results of using a Gaussian pyramid with 4 levels deteriorated, as shown in Table 4(5), possibly due to overfitting low-level features during incremental mapping. Since the image is rendered by splatting hyper primitives, the rendering speed theoretically is correlated to the number of visible hyper primitives in the current view rather than the model size of the whole scene. Although a smaller model does not necessarily imply a higher average rendering speed, a precise reconstructed model should accurately capture its essential details while using concise parameters, which is the premise for high-speed rendering. Moreover, reducing the time required for rendering enables the optimization of hyper primitives with more iterations during online mapping, ultimately leading to improved accuracy and quality.
In conclusion, the ablation study verifies that the geometry-based densification strategy allows the system to obtain sufficient hyper primitives while the Gaussian-Pyramid-based learning guarantees hyper primitives optimized thoroughly, enhancing online photorealistic mapping performance. There is no doubt that the default Photo-SLAM can reconstruct a map with more appropriate hyper primitives, and achieve better rendering quality and high rendering speed than ablated systems.
5 Conclusion
In this paper, we have proposed a novel SLAM framework called Photo-SLAM for simultaneous localization and photorealistic mapping. Instead of highly relying on resource-intensive implicit representations and neural solvers, we introduced a hyper primitives map. It enables our system to leverage explicit geometric features for localization and implicitly capture the texture information of the scenes. In addition to geometry-based densification, we proposed Gaussian-Pyramid-based learning, a new progressive training method, to further enhance mapping performance. Extensive experiments have demonstrated that Photo-SLAM significantly outperforms existing SOTA SLAMs for online photorealistic mapping. Furthermore, our system verifies its practicality by achieving real-time performance on an embedded platform, highlighting its potential for advanced robotics applications in real-world scenarios.

### 08 Apr 2024 v2

@@
 Moreover, the Photo-SLAM can run at real-time speed using an embedded platform such as Jetson AGX Orin, showing the potential of robotics applications.
+Project Page and code: https://huajianup.github.io/research/Photo-SLAM/.
 1 Introduction
-Figure 1: Photo-SLAM is a novel framework for simultaneous localization and photorealistic mapping in real-time supporting monocular, stereo, and RGB-D cameras.
-It can reconstruct high-fidelity views of scenes while render speed is up to 1000 FPS.
+Figure 1: Rendering and trajectory results. Photo-SLAM can reconstruct high-fidelity views of scenes using monocular, stereo, and RGB-D cameras while render speed is up to 1000 FPS.
 Simultaneous Localization and Mapping (SLAM) using cameras is a fundamental problem in both computer vision and robotics, seeking to enable autonomous systems to navigate and comprehend their surroundings.
@@
 Consequently, it is indispensable for them to incorporate corresponding depth information obtained from various sources such as RGB-D cameras, dense optical flow estimators [33], or monocular depth estimators [12] to ensure efficient convergence.
-Additionally, since the implicit features are decoded by the multi-layer perceptrons (MLP), it is typically necessary to carefully define a bounding area to normalize ray sampling for optimal performance [14].
+Additionally, since the implicit features are decoded by the multi-layer perceptrons (MLPs), it is typically necessary to carefully define a bounding area to normalize ray sampling for optimal performance, as discussed in [14].
 It essentially limits the scalability of the system.
@@
   • We developed the first simultaneous localization and photorealistic mapping system based on hyper primitives map.
-The novel framework supports monocular, stereo, and RGB-D cameras, in indoor and outdoor environments.
+The novel framework supports monocular, stereo, and RGB-D cameras in indoor and outdoor environments.
   • We proposed Gaussian-Pyramid-based learning allowing the model to efficiently and effectively learn multi-level features realizing high-fidelity mapping.
-  • The system, fully implemented in C++ and CUDA, achieves start-of-the-art performance and can run at real-time speed even on embedded platforms. The code will be publicly available.
+  • The system, fully implemented in C++ and CUDA, achieves start-of-the-art performance and can run at real-time speed even on embedded platforms.
 2 Related Work
@@
 The optimization in terms of position $\mathbf{P}$, rotation $\mathbf{r}$, scaling $\mathbf{s}$, density $\sigma$, and spherical harmonic coefficients $\mathbf{SH}$ is performed by minimizing the photometric loss $\mathcal{L}$ between rendering image ${I_{\text{r}}}$ and ground truth image $I_{\text{gt}}$, denoted as
-$\mathcal{L}=(1-\lambda)\left|{I_{\text{r}}-I_{\text{gt}}}\right|_{1}+\lambda\frac{1-\text{SSIM}(I_{\text{r}},I_{\text{gt}})}{2},$ (4)
+$\mathcal{L}=(1-\lambda)\left|{I_{\text{r}}-I_{\text{gt}}}\right|_{1}+\lambda{(1-\text{SSIM}(I_{\text{r}},I_{\text{gt}}))},$ (4)
 where $\text{SSIM}(I_{\text{r}},I_{\text{gt}})$ denotes structural similarity between two images and $\lambda$ is a weight factor for balance.
@@
 On Replica Dataset Localization (cm) Mapping Resources
-Cam Method RMSE $\downarrow$ STD $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Tracking FPS $\uparrow$ Rendering FPS $\uparrow$ GPU Memory Usage $\downarrow$
-Mono ORB-SLAM3 [2] 3.942 3.115 - - - 58.749 - 0
-DROID-SLAM [33] 0.725 0.308 - - - 35.473 - 11 GB
-Nice-SLAM* [45] 99.9415 35.336 16.311 0.720 0.439 2.384 0.944 12 GB
-Orbeez-SLAM [4] - - 23.246 0.790 0.336 49.200 1.030 6 GB
-Go-SLAM [43] 71.054 24.593 21.172 0.703 0.421 25.366 0.821 22 GB
-Ours (Jetson) 1.235 0.756 29.284 0.883 0.139 18.315 95.057 4 GB
-Ours (Laptop) 0.713 0.524 33.049 0.926 0.086 19.974 353.504 4 GB
-Ours 1.091 0.892 33.302 0.926 0.078 41.646 911.262 6 GB
-RGB-D ORB-SLAM3 [2] 1.833 1.478 - - - 52.209 - 0
-DROID-SLAM [33] 0.634 0.248 - - - 36.452 - 11 GB
-BundleFusion [6] 1.606 0.969 23.839 0.822 0.197 8.630 - 5 GB
-Nice-SLAM [45] 2.350 1.590 26.158 0.832 0.232 2.331 0.611 12 GB
-Orbeez-SLAM [4] 0.888 0.562 32.516 0.916 0.112 41.333 1.401 6 GB
-ESLAM [16] 0.568 0.274 30.594 0.866 0.162 6.687 2.626 21 GB
-Co-SLAM [35] 1.158 0.602 30.246 0.864 0.175 14.575 3.745 4 GB
-Go-SLAM [43] 0.571 0.218 24.158 0.766 0.352 19.437 0.444 24 GB
-Ours (Jetson) 0.581 0.289 31.978 0.916 0.101 17.926 116.395 4 GB
-Ours (Laptop) 0.590 0.289 34.853 0.944 0.062 20.597 396.082 4 GB
-Ours 0.604 0.298 34.958 0.942 0.059 42.485 1084.017 5 GB
+Cam Method RMSE $\downarrow$ STD $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Operation Time $\downarrow$ Tracking FPS $\uparrow$ Rendering FPS $\uparrow$ GPU Memory Usage $\downarrow$
+Mono ORB-SLAM3 [2] 3.942 3.115 - - - $<$1 mins 58.749 - 0
+DROID-SLAM [34] 0.725 0.308 - - - $<$2 mins 35.473 - 11 GB
+Nice-SLAM* [46] 99.9415 35.336 16.311 0.720 0.439 $>$10 mins 2.384 0.944 12 GB
+Orbeez-SLAM [4] - - 23.246 0.790 0.336 $<$5 mins 49.200 1.030 6 GB
+Go-SLAM [44] 71.054 24.593 21.172 0.703 0.421 $<$5 mins 25.366 0.821 22 GB
+Ours (Jetson) 1.235 0.756 29.284 0.883 0.139 $<$5 mins 18.315 95.057 4 GB
+Ours (Laptop) 0.713 0.524 33.049 0.926 0.086 $<$5 mins 19.974 353.504 4 GB
+Ours 1.091 0.892 33.302 0.926 0.078 $<$2 mins 41.646 911.262 6 GB
+RGB-D ORB-SLAM3 [2] 1.833 1.478 - - - $<$1 mins 52.209 - 0
+DROID-SLAM [34] 0.634 0.248 - - - $<$2 mins 36.452 - 11 GB
+BundleFusion [6] 1.606 0.969 23.839 0.822 0.197 $<$5 mins 8.630 - 5 GB
+Nice-SLAM [46] 2.350 1.590 26.158 0.832 0.232 $>$10 mins 2.331 0.611 12 GB
+Orbeez-SLAM [4] 0.888 0.562 32.516 0.916 0.112 $<$5 mins 41.333 1.401 6 GB
+ESLAM [16] 0.568 0.274 30.594 0.866 0.162 $<$5 mins 6.687 2.626 21 GB
+Co-SLAM [36] 1.158 0.602 30.246 0.864 0.175 $<$5 mins 14.575 3.745 4 GB
+Go-SLAM [44] 0.571 0.218 24.158 0.766 0.352 $<$5 mins 19.437 0.444 24 GB
+Point-SLAM [27] 0.596 0.249 34.632 0.927 0.083 $>$2 hrs 0.345 0.510 24 GB
+Ours (Jetson) 0.581 0.289 31.978 0.916 0.101 $<$5 mins 17.926 116.395 4 GB
+Ours (Laptop) 0.590 0.289 34.853 0.944 0.062 $<$5 mins 20.597 396.082 4 GB
+Ours 0.604 0.298 34.958 0.942 0.059 $<$2 mins 42.485 1084.017 5 GB
 Table 1: Quantitative results on the Replica dataset.
@@
 Zoom in for better views.
-Figure 7: Mapping comparison with other monocular camera systems.
+Figure 7: Mapping comparison with other monocular camera systems on Replica office3 scene.
 On TUM Dataset fr1-desk fr2-xyz fr3-office
@@
 Table 2: Quantitative results on the TUM RGB-D dataset. We mark the best two results with first and second.
-Refer to captionRefer to captionRefer to caption
-Refer to captionRefer to captionRefer to caption
-Refer to captionRefer to captionRefer to caption
-Figure 8: Qualitative results of Photo-SLAM on dataset TUM.
 4.1 Implementation and Experiment Setup
@@
 The compared baseline includes a SOTA classical SLAM system ORB-SLAM3 [2], a real-time RGB-D dense reconstruction system BundleFusion [6], a deep-learning based system DROID-SLAM [34], and recent SLAM systems supporting view synthesis, i.e.
-Nice-SLAM [45], Orbeez-SLAM [4], ESLAM [16], Co-SLAM [35], and Go-SLAM [43].
+Nice-SLAM [46], Orbeez-SLAM [4], ESLAM [16], Co-SLAM [36], and Point-SLAM [27] and Go-SLAM [44].
 Hardware.
@@
 Please refer to the supplementary for details.
-On Euroc Stereo ORB- SLAM3 DROID- SLAM Ours (Jetson) Ours (Laptop) Ours
-MH-01 RMSE (cm) $\downarrow$ 4.379 39.514 4.207 4.049 4.109
-PSNR $\uparrow$ - - 13.979 13.962 13.952
-SSIM $\uparrow$ - - 0.426 0.421 0.420
-LPIPS $\downarrow$ - - 0.428 0.378 0.366
-MH-02 RMSE (cm) $\downarrow$ 4.525 39.265 4.193 4.731 4.441
-PSNR $\uparrow$ - - 14.210 14.254 14.201
-SSIM $\uparrow$ - - 0.436 0.436 0.430
-LPIPS $\downarrow$ - - 0.447 0.373 0.356
-V1-01 RMSE (cm) $\downarrow$ 8.940 21.646 8.830 8.836 8.821
-PSNR $\uparrow$ - - 16.933 17.025 17.069
-SSIM $\uparrow$ - - 0.626 0.622 0.618
-LPIPS $\downarrow$ - - 0.321 0.284 0.266
-V2-01 RMSE (cm) $\downarrow$ 26.904 15.344 26.643 26.736 26.609
-PSNR $\uparrow$ - - 16.038 16.052 15.677
-SSIM $\uparrow$ - - 0.643 0.635 0.622
-LPIPS $\downarrow$ - - 0.347 0.314 0.323
-Table 3: Quantitative results on the EuRoC MAV dataset, using stereo inputs. Our Photo-SLAM is the first system to support online photorealistic mapping with stereo cameras.
-Figure 9: Mapping results of Photo-SLAM using a hand-held stereo camera in an outdoor unbounding scene.
 4.2 Results and Evaluation
@@
 9.
+Refer to captionRefer to captionRefer to caption
+Refer to captionRefer to captionRefer to caption
+Refer to captionRefer to captionRefer to caption
+Figure 8: Qualitative results of Photo-SLAM on dataset TUM.
+On Euroc Stereo ORB- SLAM3 DROID- SLAM Ours (Jetson) Ours (Laptop) Ours
+MH-01 RMSE (cm) $\downarrow$ 4.379 39.514 4.207 4.049 4.109
+PSNR $\uparrow$ - - 13.979 13.962 13.952
+SSIM $\uparrow$ - - 0.426 0.421 0.420
+LPIPS $\downarrow$ - - 0.428 0.378 0.366
+MH-02 RMSE (cm) $\downarrow$ 4.525 39.265 4.193 4.731 4.441
+PSNR $\uparrow$ - - 14.210 14.254 14.201
+SSIM $\uparrow$ - - 0.436 0.436 0.430
+LPIPS $\downarrow$ - - 0.447 0.373 0.356
+V1-01 RMSE (cm) $\downarrow$ 8.940 21.646 8.830 8.836 8.821
+PSNR $\uparrow$ - - 16.933 17.025 17.069
+SSIM $\uparrow$ - - 0.626 0.622 0.618
+LPIPS $\downarrow$ - - 0.321 0.284 0.266
+V2-01 RMSE (cm) $\downarrow$ 26.904 15.344 26.643 26.736 26.609
+PSNR $\uparrow$ - - 16.038 16.052 15.677
+SSIM $\uparrow$ - - 0.643 0.635 0.622
+LPIPS $\downarrow$ - - 0.347 0.314 0.323
+Table 3: Quantitative results on the EuRoC MAV dataset, using stereo inputs. Our Photo-SLAM is the first system to support online photorealistic mapping with stereo cameras.
+Figure 9: Mapping results of Photo-SLAM using a hand-held stereo camera in an outdoor unbounding scene.
 On Replica Mono RGB-D
@@
 Table 4: Ablation study on the effect of geometry-based densification (Geo) and Gaussian-Pyramid-based (GP) learning.
-Figure 10: Mapping results of different ablated systems on the monocular Replica scene.
 4.3 Ablation Study
@@
 The quantitative results are demonstrated in Table 4.
+Figure 10: Mapping results of different ablated systems on the monocular Replica scene.
 The relationship is tangled between rendering quality or speed, and the number of hyper primitives represented by the model size.
@@
 Furthermore, our system verifies its practicality by achieving real-time performance on an embedded platform, highlighting its potential for advanced robotics applications in real-world scenarios.
+Acknowledgements: This work is partially supported by the Innovation and Technology Support Programme of the Innovation and Technology Fund (Ref: ITS/200/20FP), the Marine Conservation Enhancement Fund (MCEF20107 & MCEF23EG01), and an internal grant from HKUST (R9429).
+Supplementary Material
+Photo-SLAM is a novel system for simultaneous localization and photorealistic mapping, which can even run on embedded platforms at real-time speed, as demonstrated in Fig.
+11.
+In this supplementary, we provide additional results regarding localization and mapping performance.
+Figure 11: The proposed system has real-time performance on embedded platforms, such as Jetson AGX Orin Developer Kit.
+Figure 12: Tracking speed comparisons on scene office0 using an RGB-D camera.
+The vertical axis denotes the processing time of each frame while the horizontal axis denotes the frame number.
+[Ave/Min/Max/Std] represent the average, minimum, and maximum tracking time and its standard deviation respectively.
+6 Localication
+Stability.
+As online systems, SLAMs are required to process the incoming frames and estimate current camera poses in time.
+Therefore, tracking stability regarding latency and the average processing time is an important factor in evaluating system performance in addition to pose estimation accuracy.
+As reported in Table 1 of the main paper, Photo-SLAM is capable of processing more than 40 frames per second with accurate pose estimation.
+The average tracking speed is about six times faster than ESLAM [16] and three times faster than Co-SLAM [36].
+Here, we provide additional analysis on tracking stability while an example plotted in Fig.
+12.
+Although the average tracking time of Go-SLAM [44] is less than Co-SLAM and ESLAM, the processing latency is high due to frequently conducting expensive global optimization.
+As shown in Fig.
+12, Go-SLAM often takes about 1 second to process the frame and estimate the pose.
+Moreover, both Nice-SLAM [46] and Co-SLAM need a longer time to accurately initialize the tracking.
+Obviously, our system can rapidly and stably process the incoming frames, having minimum average tracking time and standard deviation.
+The peak processing time of our system occurs when loop closure is detected for correcting pose estimation drift.
+Accuracy. Some qualitative tracking results of Photo-SLAM are demonstrated in Fig. 13.
+Figure 13: Trajectory in the reconstructed map. Green points denote ground truth trajectory while red denotes the estimated trajectory of Photo-SLAM.
+7 Discussion
+Online Mapping vs Offline Mapping.
+For online mapping, the mapping process occurs simultaneously with the localization process.
+Therefore it requires continuous and prompt updates with each new observation as the robot or camera moves and observes its surroundings.
+In general, online photorealistic mapping is more challenging than offline photorealistic mapping, since it is crucial to balance the trade-off between computational efficiency and rendering quality.
+As mentioned in the main paper, we proposed a geometry-based densification strategy and a Gaussian-Pyramid-based (GP) learning method to achieve high-quality online mapping.
+To further support this statement, we compared the photorealistic mapping performance between our Photo-SLAM and 3D Gaussian splatting (3DG) [18].
+3DG is the SOTA offline method which takes a set of images with known poses and a sparse point cloud as input to learn a radiance field for view synthesis.
+During the experiments, 3DG used the keyframe poses estimated by Photo-SLAM and performed training for the same duration as Photo-SLAM.
+The required point cloud input is initialized in three different ways: 1) randomly initializing 100 points; 2) randomly initializing 10,000 points; and 3) initializing from the hyper primitives map of Photo-SLAM.
+The results are reported in Table 5.
+Without inputting fine-grained point clouds, 3DG needs more time for optimization such that the rendering quality decreases.
+In addition, to enhance rendering quality, 3DG tends to densify point clouds leading to larger model size and slower rendering speed.
+Whether using monocular cameras or RGB-D cameras, Photo-SLAM consistently delivers compelling rendering quality and faster rendering speeds, owing to the effectiveness of the proposed algorithms.
+method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Rendering FPS $\uparrow$ Model Size (MB)
+1) 3DG 27.844 0.861 0.213 745.480 36.141
+2) 3DG 34.555 0.942 0.065 483.904 144.196
+3) 3DG 37.055 0.962 0.032 448.109 219.470
+Ours using Mono 33.302 0.926 0.078 911.262 31.419
+Ours using RGB-D 34.958 0.942 0.059 1084.017 35.211
+Table 5: Comparison of mapping performance between 3D Gaussian splatting (3DG) and our system Photo-SLAM with different settings on the Replica dataset.
+On TUM Dataset Resources
+Scene Cam Method Tracking FPS $\uparrow$ Rendering FPS $\uparrow$ Model Size (MB)
+fr1-desk Mono Ours (Jetson) 28.267 340.507 4.610
+Ours (Laptop) 28.330 1105.062 7.421
+Ours 57.781 2016.690 10.027
+RGB-D Ours (Jetson) 27.970 380.622 5.743
+Ours (Laptop) 28.930 1061.040 8.432
+Ours 58.378 2083.896 9.963
+fr2-xyz Mono Ours (Jetson) 24.005 169.321 14.286
+Ours (Laptop) 24.922 619.554 16.102
+Ours 58.241 1405.797 20.380
+RGB-D Ours (Jetson) 21.032 274.718 6.319
+Ours (Laptop) 22.665 701.590 13.850
+Ours 52.904 1790.120 21.399
+fr3-office Mono Ours (Jetson) 36.700 291.398 10.669
+Ours (Laptop) 38.929 824.658 16.249
+Ours 81.575 1522.120 19.211
+RGB-D Ours (Jetson) 18.039 291.907 12.726
+Ours (Laptop) 19.636 764.342 15.349
+Ours 43.650 1540.757 17.009
+Table 6: Additional results of Photo-SLAM with different platforms on the TUM dataset.
+On EuRoC Dataset Resources
+Scene Method Tracking FPS $\uparrow$ Rendering FPS $\uparrow$ Model Size (MB)
+MH-01 Ours (Jetson) 21.359 93.762 43.385
+Ours (Laptop) 25.019 316.403 89.700
+Ours 44.977 613.958 123.528
+MH-02 Ours (Jetson) 22.355 101.021 36.263
+Ours (Laptop) 26.189 332.174 81.569
+Ours 46.556 675.508 113.116
+V1-01 Ours (Jetson) 21.332 106.008 28.444
+Ours (Laptop) 25.403 367.903 55.263
+Ours 44.763 835.119 74.457
+V2-01 Ours (Jetson) 23.872 99.988 27.840
+Ours (Laptop) 27.556 307.025 62.588
+Ours 48.911 595.234 82.600
+Table 7: Additional results of Photo-SLAM with different platforms on the EuRoC MAV stereo dataset.
+On Replica Dataset Localization Mapping Resources
+[…]
+office0 Mono Ours (Jetson) 0.467 0.00334 34.415 0.940 0.949 18.328 123.551 17.703
+Ours (Laptop) 0.587 0.00343 36.227 0.954 0.071 20.422 413.645 21.703
+Ours 0.575 0.00369 36.989 0.955 0.061 42.487 930.598 24.975
+RGB-D Ours (Jetson) 0.499 0.00356 35.447 0.949 0.086 19.076 154.087 18.281
+Ours (Laptop) 0.519 0.00317 38.219 0.965 0.053 22.446 497.917 18.826
+Ours 0.522 0.00307 38.477 0.964 0.050 48.588 1447.887 19.740
+office1 Mono Ours (Jetson) 5.586 0.31140 32.382 0.904 0.113 18.312 80.048 22.904
+Ours (Laptop) 0.379 0.00463 37.970 0.954 0.060 19.968 322.160 21.224
+Ours 0.315 0.00383 37.592 0.950 0.062 42.296 857.324 26.982
+RGB-D Ours (Jetson) 0.402 0.00517 37.510 0.953 0.065 19.194 118.584 20.656
+Ours (Laptop) 0.440 0.00543 39.109 0.962 0.049 22.349 496.644 19.548
+Ours 0.436 0.00477 39.089 0.961 0.047 47.333 1263.343 21.193
+office2 Mono Ours (Jetson) 1.402 0.01452 28.083 0.900 0.131 17.502 90.181 19.704
+Ours (Laptop) 2.087 0.02154 31.202 0.927 0.098 18.975 343.662 27.332
+Ours 5.031 0.04696 31.794 0.929 0.091 39.604 930.777 31.558
+RGB-D Ours (Jetson) 1.209 0.00964 29.755 0.919 0.110 17.860 124.420 27.560
+Ours (Laptop) 1.188 0.00972 32.720 0.940 0.080 21.507 425.452 31.711
+Ours 1.276 0.01094 33.034 0.938 0.077 44.062 904.249 34.065
+office3 Mono Ours (Jetson) 0.429 0.00232 28.058 0.886 0.132 17.881 96.872 15.505
+Ours (Laptop) 0.409 0.00239 32.012 0.924 0.090 19.518 368.530 20.475
+Ours 0.472 0.00227 31.622 0.920 0.086 40.870 1131.957 26.653
+RGB-D Ours (Jetson) 0.718 0.00222 30.954 0.917 0.103 17.889 120.118 20.270
+Ours (Laptop) 0.747 0.00233 33.594 0.939 0.072 20.051 388.624 23.617
+Ours 0.782 0.00233 33.789 0.938 0.066 40.603 1125.175 25.226
+office4 Mono Ours (Jetson) 0.579 0.00305 30.399 0.921 0.109 18.755 102.949 15.201
+Ours (Laptop) 0.616 0.00279 33.656 0.940 0.078 20.311 375.033 21.444
+Ours 0.583 0.00272 34.168 0.941 0.072 42.262 849.305 26.154
+RGB-D Ours (Jetson) 0.661 0.00367 32.219 0.931 0.091 17.107 92.237 32.405
+Ours (Laptop) 0.629 0.00446 35.534 0.951 0.059 19.361 333.874 32.270
+Ours 0.582 0.00423 36.020 0.952 0.054 39.870 1061.749 35.421
+room0 Mono Ours (Jetson) 0.369 0.00321 26.423 0.787 0.221 17.987 87.246 17.121
+Ours (Laptop) 0.349 0.00294 29.899 0.868 0.125 19.521 332.127 33.151
+Ours 0.345 0.00299 29.772 0.871 0.106 41.020 754.729 44.333
+RGB-D Ours (Jetson) 0.514 0.00265 27.867 0.833 0.165 17.424 104.248 31.196
+Ours (Laptop) 0.521 0.00257 31.288 0.914 0.075 19.119 322.585 52.266
+Ours 0.541 0.00270 30.716 0.899 0.075 39.825 897.870 55.397
+room1 Mono Ours (Jetson) 0.803 0.00670 27.076 0.841 0.177 19.834 99.038 19.699
+Ours (Laptop) 1.046 0.00868 30.459 0.902 0.092 21.580 333.430 32.959
+Ours 1.183 0.00772 31.302 0.910 0.083 44.316 782.326 43.865
+RGB-D Ours (Jetson) 0.381 0.00299 30.191 0.895 0.108 18.881 121.986 29.503
+Ours (Laptop) 0.399 0.00277 33.071 0.931 0.062 21.782 367.455 43.568
+Ours 0.394 0.00320 33.511 0.934 0.057 43.352 1018.111 49.617
+room2 Mono Ours (Jetson) 0.241 0.00280 27.432 0.889 0.138 17.918 80.568 16.303
+Ours (Laptop) 0.235 0.00263 32.970 0.935 0.075 19.499 339.442 22.790
+Ours 0.225 0.00258 33.181 0.934 0.067 40.313 1053.078 26.834
+RGB-D Ours (Jetson) 0.260 0.00257 31.883 0.928 0.078 15.982 95.480 33.592
+Ours (Laptop) 0.275 0.00250 35.295 0.954 0.045 18.158 336.103 38.307
+Ours 0.305 0.00257 35.028 0.951 0.043 36.244 953.755 41.032
+Table 8: Detailed results of Photo-SLAM with different platforms on the Replica dataset.
+8 More Results
+The results of each scene of the replica dataset are detailed in Table 8 Additional qualitative results on the TUM dataset are demonstrated on Fig.
+14 and Fig.
+15, while Fig.
+16 illustrates qualitative results of Photo-SLAM on EuReC Stereo Dataset.
+Figure 14: Qualitative results of Photo-SLAM on TUM using monocular cameras.
+Figure 15: Qualitative results of Photo-SLAM on TUM using RGB-D cameras.
+Figure 16: Qualitative results of Photo-SLAM on stereo EuRoC.